# QUEST-A: Untrained Filtering with Trained Focusing led to Enhanced Quantum Architectures

### v1

@@ 6 Conclusion
In this work, we propose QUEST-A, a QAS framework that unifies discrete pruning and continuous optimization processes. This unification is achieved by decomposing the QAS problem into two subproblems: optimal circuit structure retrieval and parameter optimization. The framework comprises three core components: rapid architecture pruning mechanism, focused architecture optimization strategy, and evolutionary algorithm framework. Extensive experiments demonstrate QUEST-A’s superior performance across signal representation, image classification, and VQE tasks, validating its enhanced expressivity and precision.
Our exploration reveals that QAS requires a synergy of discrete search and continuous optimization paradigms, while leveraging validated ansatze as prior knowledge. The essence of QAS lies in its knowledge-driven dynamic nature, where the process accumulates understanding of advantageous features during circuit discovery, analogous to knowledge distillation and transfer learning in modern deep learning.
Future work will build upon these understandings of the QAS problem to design more efficient and precise search methods from the perspective of quantum circuit expressibility.
@@
Lian-Hui Yu: Methodology, Software. Xiao-Yu Li: Conceptualization, Project administration. Geng Chen: Writing - Original Draft, Visualization. Qin-Sheng Zhu: Validation. Hui Li: Formal analysis. Guo-Wu Yang: Supervision.
@@ Appendix A Preliminaries of Quantum Circuit
In QAS task, basic gates are the fundamental units of circuit design. The main task of QAS is to determine the position and parameters of these basic gates to construct the optimal quantum circuit. The basic gates used in this paper include $R_{X}$, $R_{Y}$, $R_{Z}$, and $CNOT$ gates, whose mathematical forms are as follows
$RX(\theta)=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix},RY(\theta)=\begin{pmatrix}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{pmatrix}$ (11)
@@
where, CNOT is a two-qubit gate, while the others act on single qubits. As introduced in Section  2.1, once the quantum circuit structure is determined, the variational quantum circuit can be trained through a fixed process.
It is worth noting that whether to use classical gradient optimization methods to guide PQC learning remains an open question. For instance, the promising physical neural networks (PNN) Momeni et al. (2023) approach adaptively circumvents gradient-based optimization processes.
@@ Appendix C Ablation Analysis
We conducted a set of ablation experiments using the Fashion-MNIST 10-class classification task as the experimental scenario. The experiments investigated the following three cases: (i) Training random circuits. (ii) Training after DAG screening only; (iii) Training after both DAG and KL screening;
For random circuit training (case i), we used $5$ randomly generated circuits for training. For the rapid screening cases (cases ii and iii), we screened from $100$ candidate circuits and selected the top $5$ optimal circuits. All experiments were terminated upon convergence of the loss function.
@@
Circuits screened by both DAG and KL methods (case iii) exhibited faster convergence and superior performance compared to the other two cases. Random circuits (case i) demonstrated slower convergence and inferior final performance relative to the screened circuits. Comparing random circuits with those that underwent rapid screening, it is evident that the screening process effectively altered the model’s initial performance and convergence capability, indicating the selection of circuits with higher performance potential. Moreover, the more refined screening process (case iii versus case ii) yielded further improvements in both convergence speed and final performance.

### v2

@@ 6 Conclusion
In this work, we propose QUEST-A, a knowledge-driven framework for QAS through alternating phases of untrained filtering and trained focusing. This approach views QAS as an item retrieval task with dynamic knowledge accumulation, decomposing the search process into coarse-grained retrieval guided by circuit inherent properties and fine-grained retrieval guided by task performance. Experiments demonstrate QUEST-A’s performance across signal representation, image classification, and VQE tasks, validating the effectiveness of progressive knowledge learning in architecture searching.
Our exploration reveals three key insights into quantum architecture search. First, QAS is essentially a knowledge accumulation process that starts with limited prior understanding and discovers optimal architectures through continuously enriching the knowledge of circuit characteristics. Second, validated circuit templates serve as crucial prior knowledge that determines the initial scope and efficiency of the search process. Third, effective QAS requires dynamic knowledge-driven strategies rather than static search methods, where the process progressively learns and transfers architectural knowledge, analogous to knowledge distillation in modern deep learning.
Future work should focus on developing knowledge-driven search strategies in QAS, particularly exploring the design of quantum-principle-guided circuit templates as efficient search starting points.
@@
Lian-Hui Yu: Conceptualization, Data curation, Methodology. Xiao-Yu Li: Methodology, Project administration. Geng Chen: Formal analysis, Investigation, Writing – original draft. Qin-Sheng Zhu: Formal analysis, Validation. Hui Li: Methodology, Software. Guo-Wu Yang: Software.
@@ Appendix A Preliminaries of Quantum Circuit
In QAS task, basic gates are the fundamental units of circuit design. The core challenge lies in discovering optimal architectures through progressively understanding the characteristics and combinations of these basic gates. The basic gates used in this paper include $R_{X}$, $R_{Y}$, $R_{Z}$, and $CNOT$ gates, whose mathematical forms are as follows:
$RX(\theta)=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix},RY(\theta)=\begin{pmatrix}\cos(\theta/2)&-\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{pmatrix},$ (11)
@@
where, CNOT is a two-qubit gate, while the others act on single qubits. Similar to NAS, the challenge of QAS stems from the exponentially growing search space due to gate combinations and task-specific search objectives.
@@ Appendix C Ablation Analysis
We conducted a set of ablation experiments to validate the effectiveness of progressive knowledge learning using the Fashion-MNIST 10-class classification task as the experimental scenario. The experiments investigated the following three cases: (i) Trained focusing in random circuits . (ii) Trained focusing after DAG-based untrained filtering only; (iii) Trained focusing after DAG and KL-based untrained filtering;
For random trained focusing (case i), we used $5$ randomly generated circuits for focusing. For the untrained filtering cases (cases ii and iii), we screened from $100$ candidate circuits and selected the top $5$ optimal circuits. All experiments were terminated upon convergence of the loss function.
@@
Case (iii) with complete untrained filtering exhibited faster convergence and superior performance compared to the other two cases. Case (i) with random circuits demonstrated slower convergence and inferior final performance. Comparing case (i) with cases using untrained filtering, it is evident that the filtering process effectively enhanced the initial search space and convergence capability, indicating better architecture discovery potential. Moreover, the comprehensive untrained filtering process in case (iii) versus case (ii) yielded further improvements in both convergence speed and final performance.